Tom,Bob and Joe are brothers. their combined ages are 47. Joe was born 3 years before Bob and 4 years after Tom. How old is each brother?
Answer by rfer(16322)

tom x=14
bob x+4=18
joe x+1=15
x+x+4+x+1=47
3x+5=47
3x=42
x=14
x+4=18
x+1=15